# Dynamics of contaminant flow through porous media containing biochar adsorbers

### v1

@@ 5 Conclusion
If one extends these results to application recommendations, we can state that the observed differences between the mono- and polydisperse cases are larger than those observed when biochar is present or not. This would correspond to cases 1 and 2 (no biochar) and 5 and 6 (biochar with a limited capacity). The concept of limited adsorption capacity is analogous to the realistic adsorbance occurring in the experimental biochar, which cannot remove the solute indefinitely. These cases only slightly outperform the cases in which there are no reactive particles present. This is due to the rate at which the limited capacity particles reach capacity and become inert. Cases 3 and 4 have unlimited capacity, which one could say is analogous with roots or some other living biological matter that does not have a capacity in the same manner as biochar. The practical implication here is that monodisperse beds are preferred as they promote more even distribution of the contaminant within the soil and thus allows the biochar to act with maximum effect. Additionally, the presence of biological matter with an unlimited adsorption capacity must be taken into account as this can dramatically effect the results, even with an adsorption rate the same as any biochar present. This effect will obviously increase if the adsorption rate is higher.
If one wishes to increase the efficacy of biochar, a high absorptive rate with a similarly high capacity will greatly reduce breakthrough. This can be accomplished by modifying the material itself and/or increasing the total quantity of biochar within the soil, which will consequently increase its capacity. Current practice suggests no more than 15% biochar by volume is beneficial given common rainfall quantities; however our results show that performance can be increased if one takes into account the factors mentioned above, with the additional mention of total reactive surface area and the idea of induced preferential pathing. Lower rainfall intensities will cause higher residence time of the concentration front near the biochar; however once they saturate, which we show can take place quite quickly, breakthrough will still occur.
Thus a more accurate method for modeling the capacity of different adsorbers should be investigated to improve the predictive ability of the models. Finally, given that monodisperse particles promote a more homogeneous concentration front and that total reactive area matters more than individual particle size we suggest that more uniform sized substrates are used.

### v2

@@ 5 Conclusion
We have chosen to use spherical packed beds over a realistic soil substrate geometry due to the intrinsic inhomogeneity found within soil scans taken using x-ray microtomography (XMT) or any similar technique. A chosen domain may contain few very large particles and multitudes of tiny particles or anywhere in between, rendering the porosity of any such domains fluctuating wildly, making comparisons difficult. In future it is worthwhile to determine if the trends observed in this work hold true for the realistic soils as well.
If one extends these results to application recommendations, we can state that the observed differences between the mono- and polydisperse cases are larger than those observed when biochar is present or not. This would correspond to cases 1 and 2 (no biochar) and 5 and 6 (biochar with a limited capacity). The concept of limited adsorption capacity is analogous to the realistic adsorbance occurring in the experimental biochar, which cannot remove the solute indefinitely. These cases only slightly outperform the cases in which there are no reactive particles present. This is due to the rate at which the limited capacity particles reach capacity and become inert. Cases 3 and 4 have unlimited capacity, which one could say is analogous with roots or some other living biological matter that does not have a capacity in the same manner as biochar.
The practical implication here is that monodisperse beds are preferred as they promote more even distribution of the solute within the domain and thus allow the random adsorbers to act with maximum effect, i.e. reaction-limited. Additionally, the presence of adsorbers with an unlimited adsorption capacity must be taken into account as this can dramatically effect the results, particularly in proximity to each other. This effect will obviously increase if the adsorption rate is increased.
If one wishes to increase the efficacy of biochar, a high absorptive rate with a similarly high capacity will greatly reduce breakthrough. This can be accomplished by modifying the material itself and/or increasing the total quantity of biochar within the soil, which will consequently increase its capacity. Current practice suggests no more than 15% biochar by volume is beneficial given common rainfall quantities; however our results show that performance can be increased if one takes into account the factors mentioned above, with the additional mention of total reactive surface area and the idea of induced preferential pathing. Lower rainfall intensities will cause higher residence time of the concentration front near the biochar; however once they saturate, which we show can take place quite quickly, breakthrough will still occur. Thus a more accurate method for modeling the capacity of different adsorbers should be investigated to improve the predictive ability of the models. Finally, given that monodisperse particles promote a more homogeneous concentration front and that total reactive area matters more than individual particle size we suggest that more uniform sized substrates are used.